Distance between second and third shell of hydrogen atom is - (1)1.058×10^-8 cm(2)0.529×10^-8 cm(3)2.116×10^-8 cm(4)2.645×10^-8 cmExplain?????

Grade:11

1 Answers

Vikas TU
7 years ago
Radii for second orbit of H – atom is:
r1 = 0.0529*2^2/1 
Radii for third shell:
r2 = 0.0529*3^2/1 
Distance b/w second and third orbit is hence => r2 – r1 => 0.0529*3^2/1  – 0.0529*2^2/1 = > 0.0529*3 = > 0.158 x 10^-10 meter.
that is close to first option.